# NuLite - Lightweight and Fast Model for Nuclei Instance Segmentation and Classification

### v1

@@ Abstract
In pathology, accurate and efficient analysis of Hematoxylin and Eosin (H&E) slides is crucial for timely and effective cancer diagnosis. Although many deep learning solutions for nuclei instance segmentation and classification exist in the literature, they often entail high computational costs and resource requirements, thus limiting their practical usage in medical applications. To address this issue, we introduce a novel convolutional neural network, NuLite, a U-Net-like architecture designed explicitly on Fast-ViT, a state-of-the-art (SOTA) lightweight CNN. We obtained three versions of our model, NuLite-S, NuLite-M, and NuLite-H, trained on the PanNuke dataset. The experimental results prove that our models equal CellViT (SOTA) in terms of panoptic quality and detection. However, our lightest model, NuLite-S, is 40 times smaller in terms of parameters and about 8 times smaller in terms of GFlops, while our heaviest model is 17 times smaller in terms of parameters and about 7 times smaller in terms of GFlops. Moreover, our model is up to about 8 times faster than CellViT. Lastly, to prove the effectiveness of our solution, we provide a robust comparison of external datasets, namely CoNseP, MoNuSeg, and GlySAC. Our model is publicly available at https://github.com/CosmoIknosLab/NuLite.
@@
journal: Arxiv
@@ 1 Introduction
Cancer is a disease concerned with the uncontrolled growth and spread of abnormal cells, a significant global health challenge [1]. Accurate diagnosis is essential in cancer treatment because it enables targeted therapies that improve patient outcomes and the chance of recovery. Advancements in computer vision techniques have significantly affected computational pathology (CPATH), opening new frontiers for analyzing histopathological images, like the Hematoxylin and Eosin (H&E) stained one [2]. The precise segmentation and classification of cells became an exciting task in the literature due to the importance of understanding the morphology and topology of tissue in cancer diagnosis [3]. However, this task in complex tissue environments poses many challenges due to the heterogeneity and overlap of nuclei structures, demanding robust and efficient solutions [4]. To address these challenges, recent research has focused on developing sophisticated algorithms that leverage deep learning techniques, demonstrating superior performance in various image analysis tasks. These algorithms are designed to accurately identify and classify cellular components, even in complex and heterogeneous tissue environments, thereby assisting pathologists in making more informed diagnostic decisions. Moreover, integrating machine learning models with domain-specific knowledge, such as the spatial relationships and morphological features of cells, has further enhanced the accuracy and robustness of computational pathology tools. This synergy between advanced computational methods and medical expertise promises to significantly advance the field of cancer diagnostics, offering the potential for more personalized and effective treatment plans.
@@
Over the years, scholars have proposed many methods to overcome traditional barriers encountered in histopathological analysis in different tasks [5, 6]. In particular, many deep learning solutions have shown promising results for nuclei instance segmentation and classification tasks, starting with the introduction of U-Net [7]. Furthermore, advanced neural network architectures, like ResNet [8] and Vision Transformer (ViT) [9], offered sophisticated mechanisms for learning detailed features and patterns without the constraints imposed by prior techniques, further improving the effectiveness of new models. The recent trend toward integrating different modalities of deep learning, such as Convolutional Neural Networks (CNNs) combined with structures like U-Nets or multi-branch networks like HoVer-Net [10], demonstrates the field’s evolution toward more precise and robust techniques. Additionally, implementing spatial and morphological constraints within network architectures further refine their output, ensuring that cell segmentation is precise and contextually appropriate. Moreover, a recent technique, CellViT [11], demonstrated using ViT to address nuclei instance segmentation and classification tasks, achieving the SOTA results.
@@
In this manuscript, we present NuLite, a new UNet-like CNN [7] architecture designed explicitly for segmenting and classifying nuclei instances in Hematoxylin and Eosin (H&E) images. Our architecture consists of the FastViT [12] encoder with three decoders, each purpose-built to perform one of the tasks identified in the HoVer-Net: nuclei prediction, horizontal and vertical map prediction, and nuclei classification [10]. NuLite is a faster, lighter alternative with state-of-the-art (SOTA) panoptic quality and detection performance. We proved its efficacy and efficiency through rigorous testing on benchmark datasets such as PanNuke [13]. Furthermore, we conducted comprehensive evaluations on additional datasets such as MoNuSeg [14], CoNSeP [10], and GlySAC [15]. NuLite consistently achieved SOTA results in these tests, outperforming advanced models like CellViT in various metrics, including precision, recall, and F1-score. These evaluations underscore the robustness and generalizability of NuLite across different types of histopathological images, highlighting its potential as a versatile tool in computational pathology.
@@ 2.1.2 Deep learning approaches
In the last decade, deep learning techniques leveraged the limitations of traditional approaches. One of the first networks that achieved promising results in nuclei segmentation, posing the basis for all modern techniques, was U-Net proposed by Olaf Ronneberger et al. [7]. U-Net is an encoder-decoder neural network with skip connections, which helps preserve details crucial for medical image analysis. However, its original version proposed a way to separate clustered nuclei, which is a significant challenge in histopathology. Another network was BRP-Net [24] that creates nuclei proposals in the first place, then refines the boundary, and finally creates a segmentation out of this. However, this approach resulted in computationally intensive and slow. Similarly, Alemi et al. introduced Mask-RCNN [25], built on Fast-RCNN [26], adding a segmentation branch after nuclei detection. Instead, Raza et al. proposed Micro-Net [27] updating U-Net to handle nuclei of varying sizes. Another network that significantly improved the nuclei instance segmentation and classification is HoVer-Net [10], which has U-Net architecture with three branches that predict nuclei against the background, vertical and horizontal map, and nuclei types. The vertical and horizontal maps are crucial to separate overlapped nuclei and, in general, to perform instance segmentation. Following the idea of [10], the authors in [11] proposed CellViT, which follows the same architecture but employs a ViT as the encoder, and the authors designed a decoder inspired by UNETR [28]. Instead, authors in [29] proposed a framework to obtain a smaller and lighter model than HoVerNet, HoVer-UNet, that is, a U-Net-like neural network with one decoder trained using a knowledge distillation approach. Other recent networks proposed in the literature are STARDIST [30] and CPP-Net [31], which used star-convex polygons for segmentation, with CPP-Net enhancing the model by integrating shape-aware loss functions to improve accuracy. Similarly, TSFD-Net [32] employed a Feature Pyramid Network and integrated a tissue-classifier branch to handle tissue-specific features, using advanced loss functions to manage class imbalance. Moreover, the SONNET [15] network is a deep learning model designed for simultaneous segmentation and classification of nuclei in large-scale multi-tissue histology images. It employs a self-guided ordinal regression approach that stratifies nuclear pixels based on their distance from the center of mass, improving the accuracy of segmenting overlapping nuclei.
@@ 3.1 NuLite architecture
We designed NuLite with a U-Net-like architecture and three decoders, utilizing FastViT [12], a state-of-the-art Vision Transformer known for its lightweight and efficient design as the backbone. The decision to use a U-Net structure is based on its proven success in medical image segmentation tasks, which aligns with our goal of accurate nuclei detection and classification. Our approach draws inspiration from HoVer-Net [10], which has established itself in the literature as an effective method for nuclei segmentation. Therefore, our model predicts nuclei maps, type maps, and horizontal and vertical maps, followed by a watershed algorithm to perform nuclei instance segmentation in a postprocessing step. Thus, our network comprises three specialized decoders: the NP-BRANCH for nuclei segmentation against the background, the HV-BRANCH for predicting horizontal and vertical orientation maps, and the NC-BRANCH for nuclei classification, as illustrated in Figure 1. This architecture supports detailed nuclei analysis through a postprocessing step that leverages the NP-BRANCH and HV-BRANCH outputs to precisely detect individual nuclei, subsequently using the NC-BRANCH for classification. We carefully designed the decoders to minimize computational overhead, maintaining low parameter counts and GFlops, thus ensuring efficiency. Furthermore, we integrated a dense layer within the encoder to facilitate tissue classification, extending the functionality of the network beyond nuclei analysis, and results useful during the training step to improve the segmentation and classification capabilities.
Figure 1: NuLite architecture. The network takes as input multiple feature maps, each encoding different spatial and contextual information. These features are then processed through convolutional and pooling layers, organized into distinct stages and branches. The network employs patch embedding to capture local patterns and leverages deconvolutional layers to generate a high-resolution segmentation map. The final output is a pixel-wise classification that assigns each pixel to a specific tissue class.
We focused on the decoder design, and we built it to work with the FastViT encoder. Therefore, the decoder consists of six main layers, as detailed in Table 1.
@@ 5 Conv2D (3x3) - BN - ReLU
6 Conv2D (3x3) - BN - ReLU
@@
As a standard U-like architecture, we designed the skip connection between the main block output of the encoder, namely stage 1 to stage 4, as shown in Figure 1, and each main block of our branches. Furthermore, we added a skip connection between the original output after a convolutional layer and the last layer of each decoder. The architecture of each decoder comprises six functional layers, as described in Table 1. These layers operate on input images defined by dimensions height $H$ and width $W$, with $Z$ indicating the number of channels from the encoder. The first, fourth, and fifth layers are structured with a $3\times 3$ convolutional layer, which is succeeded by batch normalization and ReLU activation and augmented by a deconvolution layer. The output from the first and fourth layers yields feature maps with half the number of channels of $Z$ but with dimensions expanded to twice the height ($2H$) and width ($2W$). The fifth layer, however, maintains the channel count of $Z$ and doubles the height and width. The design of the second and third layers integrates two $3\times 3$ convolutional layers, each followed by batch normalization and ReLU activation. These layers are completed by a deconvolution layer that produces outputs with a quarter of the channels of $Z$ and twice the original dimensions in height and width. The final, sixth layer includes a $3\times 3$ convolution followed by batch normalization and ReLU, culminating in a $1\times 1$ convolution that adjusts the output channels to meet specific requirements. As notable from the decoder structure, the second and third layers contain two convolutional blocks, while the rest have only one covolutional block. That is because the number of feature maps is reduced by 4 times in the second and third and 2 times in the rest. FastViT exists in several configurations, distinguished by their pecifications. The available variants include T8, T12, S12, SA12, SA24, SA36, and MA36. The parameter $Z$, which denotes the number of channels, varies across these models. Specifically, $Z$ is set to 384 for T8, while it remains consistent at 512 for T12, S12, SA12, SA24, and SA36. For the MA36 configuration, $Z$ increases to 608. Therefore, in this paper, we consider a server version of NuLite, each using a version of FastViT.
@@ 3.2 Loss fuction
To train NuLite, we use a combination of different loss functions for each network branch, as also suggested in [10, 11]. Therefore, the total loss is defined as the sum of a loss for each branch, as shown in Equation 1.
@@
$L_{NP}$ is the loss for the NP-branch, defined as a linear combination of Focal Tversky loss (FTL) and Dice loss (DICE), as shown in Equation 2.
@@
$L_{HV}$ is the loss for the HV-branch, defined as a linear combination of Mean Square Error (MSE) and Mean Square Gradient Error (MSGE), as shown in Equation 3.
@@
$L_{nt}$ is the loss for the NT-branch, defined as a linear combination of FLT, DICE, and Binary Cross Entropy Loss (BCE) as shown in Equation 4.
@@
$L_{TC}$ the loss for the TC-branch, computed as Cross Entropy (CE) as shown in Equation 5
@@
In these equations, $\lambda_{\text{brach}^{\text{loss}}}$ coefficients represent the weight given to each loss component within a branch.
@@ 3.3 Post-Processing
As described in the preview sections, our network, NuLite, encloses three specialized branches dedicated to extracting essential information for a nuclei instance segmentation and classification postprocessing step. Due to our network following the idea proposed in HoVer-Net and Cell-ViT, postprocessing is a crucial step in refining the raw predictions produced by the network. NP-BRANCH output is a probability map indicating the likelihood of each pixel belonging to a nucleus. A threshold is applied to this probability map to generate a binary mask. Pixels with probabilities above the threshold are considered part of a nucleus. HV-BRANCH contains horizontal and vertical gradient maps (HV maps) that help to delineate the nuclei boundaries more accurately. These maps provide additional information about the direction and magnitude of changes in the image, which is useful for refining the edges of the segmented nuclei. The gradient information from the HV maps is used to split merged nuclei. This process is critical when there are overlapped nuclei. Each nucleus identified from the segmentation step is classified according to the output of TP-BRANCH. Finally, some morphological operations, like dilation and erosion, can be applied to smooth the boundaries of the segmented nuclei and remove small noise artifacts, improving the visual quality of the segmentation masks.

### v2

@@ Abstract
In pathology, accurate and efficient analysis of Hematoxylin and Eosin (H&E) slides is crucial for timely and effective cancer diagnosis. For these reasons, nuclei instance segmentation and classification tools are helpful, allowing pathologists to detect and identify regions of interest and perform quantitive analysis. Although many deep-learning solutions for this task exist in the literature, they often entail high computational costs and resource requirements, thus limiting their practical usage in medical applications. To address this issue, we introduce NuLite, a U-Net-like architecture designed explicitly to be lightweight and fast. We obtained three versions of our model, NuLite-S, NuLite-M, and NuLite-H, trained on the PanNuke dataset. The experimental results prove that our models are equivalent to CellViT (SOTA) in terms of panoptic quality and F-score. However, our lightest model, NuLite-T, is about 58 times smaller in terms of parameters and about 10 times smaller in terms of GFlops. In comparison, our heaviest model is about 15 times smaller in terms of parameters and about 7 times smaller in terms of GFlops. Moreover, considering the GPU latency, our model is up to about 13 times faster than CellViT. Lastly, to prove the effectiveness of our solution, we provide a robust comparison of external datasets, namely CoNseP, MoNuSeg, and GlySAC. Our model is publicly available at https://github.com/CosmoIknosLab/NuLite.
@@
journal: Medical Image Analysis
@@ 1 Introduction
Cancer is a disease concerned with the uncontrolled growth and spread of abnormal cells, a significant global health challenge [1]. Accurate diagnosis is essential in cancer treatment because it enables targeted therapies that improve patient outcomes and the chance of recovery. Advancements in computer vision techniques have significantly affected computational pathology (CPATH), opening new frontiers for analyzing histopathological images, like the Hematoxylin and Eosin (H&E) stained one [2]. The precise segmentation and classification of cells became an exciting task in the literature due to the importance of understanding the morphology and topology of tissue in cancer diagnosis [3]. However, this task in complex tissue environments poses many challenges due to the heterogeneity and overlap of nuclei structures, demanding robust and efficient solutions [4]. To address these challenges, recent research has focused on developing sophisticated algorithms that leverage deep learning techniques, demonstrating superior performance in various image analysis tasks. These algorithms are designed to accurately identify and classify cellular components, even in complex and heterogeneous tissue environments. However, it is essential to note that these tools are meant to supplement pathologists and assist them in making more informed diagnostic decisions. Moreover, integrating machine learning models with domain-specific knowledge, such as the spatial relationships and morphological features of cells, has further enhanced the accuracy and robustness of computational pathology tools. This synergy between advanced computational methods and pathologists’ expertise promises to significantly advance the field of cancer diagnostics, offering the potential for more personalized and effective treatment plans.
@@
Over the years, scholars have proposed many methods to overcome traditional barriers encountered in histopathological analysis in different tasks [5, 6, 7]. In particular, many deep learning solutions have shown promising results for nuclei instance segmentation and classification tasks, starting with the introduction of U-Net [8]. Furthermore, advanced neural network architectures, like ResNet [9] and Vision Transformer (ViT) [10], offered sophisticated mechanisms for learning detailed features and patterns without the constraints imposed by prior techniques, further improving the effectiveness of new models. The recent trend toward integrating different modalities of deep learning, such as Convolutional Neural Networks (CNNs) combined with structures like U-Nets or multi-branch networks like HoVer-Net [11], demonstrates the field’s evolution toward more precise and robust techniques. Additionally, implementing spatial and morphological constraints within network architectures further refine their output, ensuring that cell segmentation is precise and contextually appropriate. Moreover, a recent technique, CellViT [12], demonstrated using ViT to address nuclei instance segmentation and classification tasks, achieving the SOTA results.
@@
This manuscript presents NuLite, a new UNet-like CNN [8] architecture designed for segmenting and classifying nuclei instances in Hematoxylin and Eosin (H&E) images. Our architecture consists of the FastViT [13] encoder, one decoder, and three segmentation heads purpose-built to perform one of the tasks identified in the HoVer-Net: nuclei prediction, horizontal and vertical map prediction, and nuclei classification [11]. We decided to use one decoder, contrary to what is commonly reported in the literature, to avoid parameter redundancy among the decoders and further reduce the parameters and GFLOPS. NuLite is a faster, lighter alternative with state-of-the-art (SOTA) panoptic quality and detection performance. We proved its efficacy and efficiency through rigorous testing on benchmark datasets such as PanNuke [14]. Furthermore, we conducted comprehensive evaluations on additional datasets such as MoNuSeg [15], CoNSeP [11], and GlySAC [16]. NuLite consistently achieved SOTA results in these tests, outperforming advanced models like CellViT in various metrics, including precision, recall, and F1-score. These evaluations underscore the robustness and generalizability of NuLite across different types of histopathological images, highlighting its potential as a versatile tool in computational pathology.
@@ 2.1.2 Deep learning approaches
In the last decade, deep learning techniques leveraged the limitations of traditional approaches. One of the first networks that achieved promising results in nuclei segmentation, posing the basis for all modern techniques, was U-Net proposed by Olaf Ronneberger et al. [8]. U-Net is an encoder-decoder neural network with skip connections, which helps preserve details crucial for medical image analysis. However, its original version proposed a way to separate clustered nuclei, which is a significant challenge in histopathology. Another network was BRP-Net [25] that creates nuclei proposals in the first place, then refines the boundary, and finally creates a segmentation out of this. However, this approach resulted in computationally intensive and slow. Similarly, Alemi et al. introduced Mask-RCNN [26], built on Fast-RCNN [27], adding a segmentation branch after nuclei detection. Instead, Raza et al. proposed Micro-Net [28] updating U-Net to handle nuclei of varying sizes. Another network that significantly improved the nuclei instance segmentation and classification is HoVer-Net [11], which has U-Net architecture with three branches that predict nuclei against the background, vertical and horizontal map, and nuclei types. The vertical and horizontal maps are crucial to separate overlapped nuclei and, in general, to perform instance segmentation. Following the idea of [11], the authors in [12] proposed CellViT, which follows the same architecture but employs a ViT as the encoder, and the authors designed a decoder inspired by UNETR [29]. Instead, authors in [30] proposed a framework to obtain a smaller and lighter model than HoVerNet, HoVer-UNet, that is, a U-Net-like neural network with one decoder trained using a knowledge distillation approach. Other recent networks proposed in the literature are STARDIST [31], and CPP-Net [32], which used star-convex polygons for segmentation, with CPP-Net enhancing the model by integrating shape-aware loss functions to improve accuracy. Similarly, TSFD-Net [33] employed a Feature Pyramid Network and integrated a tissue-classifier branch to handle tissue-specific features, using advanced loss functions to manage class imbalance. Moreover, the SONNET [16] network is a deep learning model designed for simultaneous segmentation and classification of nuclei in large-scale multi-tissue histology images. It employs a self-guided ordinal regression approach that stratifies nuclear pixels based on their distance from the center of mass, improving the accuracy of segmenting overlapping nuclei.
@@ 2.2 Vision Transformers
Figure 1: NuLite architecture. The network has a U-Net-like architecture with one decoder and three segmentation heads: one to predict nuclei, one to predict horizontal and vertical maps, and one to predict nuclei types. The post-processing uses all outputs to perform nuclei instance segmentation and assign the predicted nucleus type to each one.
@@ 3.1 NuLite architecture
We designed NuLite with a U-Net-like architecture and three decoders, utilizing FastViT [13], a state-of-the-art Vision Transformer known for its lightweight and efficient design as the backbone. We decided to use U-Net architecture for its proven success in medical image segmentation tasks, as also shown by other recent models such as CellViT [12] and StartDist [31], which aligns with our goal of accurate nuclei detection and classification. Our approach draws inspiration from HoVer-Net [11], which has established itself in the literature as an effective method for nuclei segmentation, as demonstrated by CellViT [12], HoVer-NeXt [38], and HoVer-UNet [30], but using just one decoder with three segmentation heads. Therefore, our model predicts nuclei maps, type maps, and horizontal and vertical maps, followed by a watershed algorithm to perform nuclei instance segmentation in a postprocessing step. Thus, our network comprises three heads: the NP-HEAD for nuclei segmentation against the background, the HV-HEAD for predicting horizontal and vertical orientation maps, and the NC-HEAD for nuclei classification, as illustrated in Figure 1. This architecture supports detailed nuclei analysis through a postprocessing step that leverages the NP-MAP and HV-MAPS to precisely detect individual nuclei, subsequently using the NC-MAP to assign the type to each nucleus instance. We carefully designed the decoders to minimize computational overhead, maintaining low parameter counts and GFlops, thus ensuring efficiency. Furthermore, we integrated a dense layer within the encoder to facilitate tissue classification, extending the functionality of the network beyond nuclei analysis, and results useful during the training step to improve the segmentation and classification capabilities.
We focused on the decoder design and built it to work with the FastViT [13] encoder. Therefore, the decoder consists of five main layers and three segmentation heads, as detailed in Table 1.
@@
NP.HEAD Conv2D (3x3) - BN - ReLU
Conv2D (1x1)
$2\cdot Z\times H\times W$ $2\times H\times W$
HV.HEAD Conv2D (3x3) - BN - ReLU
Conv2D (1x1)
$2\cdot Z\times H\times W$ $2\times H\times W$
NC.HEAD Conv2D (3x3) - BN - ReLU
@@
As a standard U-like architecture, we employ the skip connection between the main block output of the encoder, namely stage 1 to stage 4, as shown in Figure 1, and each main block of our decoder. Furthermore, we add a skip connection between the original input after a convolutional layer and the last layer of the decoder. The decoder architecture comprises five layers, namely DEC.1, DEC.2, DEC.3, DEC.4, and DEC.5, and three segmentation heads, namely NP.HEAD, HV.HEAD, and NC.HEAD, as described in Table 1. These layers operate on input images defined by dimensions height $H$ and width $W$, with $Z$ indicating the number of channels output from the encoder. We structured the DEC.1, DEC.4, and DEC.5 with a $3\times 3$ convolutional layer, which is succeeded by batch normalization and ReLU activation and augmented by a deconvolution layer. The output from DEC.1 and DEC.4 yields feature maps with half the number of channels of $Z$ but with dimensions expanded to twice the height ($2H$) and width ($2W$). However, DEC.5 maintains the channel count of $Z$ and doubles the height and width. The design of DEC.2 and DEC.3 integrates two $3\times 3$ convolutional layers, each followed by batch normalization and ReLU activation. These layers are completed by a deconvolution layer that produces outputs with a quarter of the channels of $Z$ and twice the original dimensions in height and width. Each segmentation head has a $3\times 3$ convolution followed by batch normalization and ReLU, followed by a $1\times 1$ convolution that adjusts the output channels to meet specific requirements, namely 2 channels for nuclei prediction and horizontal and vertical map prediction and $C$ channels for nuclei classification where $C$ is the number of class contained in the training dataset. As notable from the decoder structure, the second and third layers contain two convolutional blocks, while the rest have only one covolutional block. That is because the number of feature maps is reduced by 4 times in the second and third and 2 times in the rest. FastViT exists in several configurations, including T8, T12, S12, SA12, SA24, SA36, and MA36. The parameter $Z$, which denotes the number of channels, varies across these models. Specifically, $Z$ is set to 384 for T8, while it remains consistent at 512 for T12, S12, SA12, SA24, and SA36. For the MA36 configuration, $Z$ increases to 608. Therefore, in this paper, we consider a server version of NuLite, each using a version of FastViT.
@@ 3.2 Loss fuction
To train NuLite, we use a combination of different loss functions for each network output, as also suggested in [11, 12]. Therefore, the total loss is defined as the sum of a loss for each segmentation head, as shown in Equation 1.
@@
$L_{NP}$ is the loss for the NP-HEAD, defined as a linear combination of Focal Tversky loss (FTL) and Dice loss (DICE), as shown in Equation 2.
@@
$L_{HV}$ is the loss for the HV-HEAD, defined as a linear combination of Mean Square Error (MSE) and Mean Square Gradient Error (MSGE), as shown in Equation 3.
@@
$L_{nt}$ is the loss for the NT-HEAD, defined as a linear combination of FLT, DICE, and Binary Cross Entropy Loss (BCE) as shown in Equation 4.
@@
$L_{TC}$ the loss for the tissue classification, computed as Cross Entropy (CE) as shown in Equation 5
@@
In these equations, $\lambda_{\text{brach}^{\text{loss}}}$ coefficients represent the weight given to each loss component.
@@ 3.3 Post-Processing
As described in the preview sections, our network, NuLite, encloses three specialized segmentation heads dedicated to extracting essential information for a nuclei instance segmentation and classification postprocessing step. Due to our network following the idea proposed in HoVer-Net and Cell-ViT, postprocessing is a crucial step in refining the raw predictions produced by the network. NP-HEAD output is a probability map indicating the likelihood of each pixel belonging to a nucleus. A threshold is applied to this probability map to generate a binary mask. Pixels with probabilities above the threshold are considered part of a nucleus. HV-HEAD contains horizontal and vertical gradient maps (HV maps) that help to delineate the nuclei boundaries more accurately. These maps provide additional information about the direction and magnitude of changes in the image, which is useful for refining the edges of the segmented nuclei. The gradient information from the HV maps is used to split merged nuclei. This process is critical when there are overlapped nuclei. Each nucleus identified from the segmentation step is classified according to the output of NC-HEAD. Finally, some morphological operations, like dilation and erosion, can be applied to smooth the boundaries of the segmented nuclei and remove small noise artifacts, improving the visual quality of the segmentation masks.